OpenStudy (magbak):

For problems 1 and 2, graph the data points. Draw the line of best fit and, using complete sentences, explain how you found it in 3-4 sentences. Write the equation for the line of best fit in standard form. 1. (10, 150) (20, 275) (30, 340) (40, 400) (50, 475) Title: Intersection Cars Horizontal axis: Minutes elapsed Vertical axis: Total cars passing through 2. (2, 12) (4, 21) (6, 19) (8, 30) (10, 29) (12, 26) (14, 35) (16, 40) (18, 46) (20, 42) Title: Savings Account Horizontal axis: Years Vertical axis: Balance in $ × 1000

thomaster (thomaster):

So what's the question?

OpenStudy (magbak):

For problems 1 and 2, graph the data points. Draw the line of best fit and, using complete sentences, explain how you found it in 3-4 sentences. Write the equation for the line of best fit in standard form. Also it is 2 separate questions.

thomaster (thomaster):

 That would be the first graph

OpenStudy (magbak):

And the line of best fit for the first graph is ......

thomaster (thomaster):

It's a linear graph so the general equation is \(\Large y=ax+b\) a = the slope b = the intercept (the point where the line crosses the y-axis you can calculate the slope by picking 2 points on the line and divide the y values by the x values in this way: \(\large\dfrac{(Y_2-Y1)}{(X_2-X_1)}\)

thomaster (thomaster):

So for example let's take point (20,275) and (40,400) The slope will be \(\large\dfrac{(400-275)}{(40-20)}=6.25\)

OpenStudy (magbak):